Find the equation of the line containing (-2,-3) with m= -3/4.

Answer by nerdybill(7384)

Find the equation of the line containing (-2,-3) with m= -3/4.
Plug the given information into the "point-slope" form:
y - y1 = m(x - x1)
y - (-3) = (-3/4)(x - (-2))
y + 3 = (-3/4)(x + 2)
y + 3 = (-3/4)x - 3/2
y = (-3/4)x - 3/2 - 3
y = (-3/4)x - 3/2 - 6/2
y = (-3/4)x - 9/2